# Ego-Vision World Model for Humanoid Contact Planning

### v1

@@ Abstract
Enabling humanoid robots to exploit physical contact, rather than simply avoid collisions, is crucial for autonomy in unstructured environments. Traditional optimization-based planners struggle with contact complexity, while on-policy reinforcement learning (RL) is sample-inefficient and has limited multi-task ability. We propose a framework combining a learned world model with sampling-based Model Predictive Control (MPC), trained on a demonstration-free offline dataset to predict future outcomes in a compressed latent space. To address sparse contact rewards and sensor noise, the MPC uses a learned surrogate value function for dense, robust planning. Our single, scalable model supports contact-aware tasks, including wall support after perturbation, blocking incoming objects, and traversing height-limited arches, with improved data efficiency and multi-task capability over on-policy RL. Deployed on a physical humanoid, our system achieves robust, real-time contact planning from proprioception and ego-centric depth images. Website: https://ego-vcp.github.io/
@@ I Introduction
Humanoids are expected to advance from dynamic locomotion [1, 2] to intelligent interaction in complex, unstructured environments. Achieving this requires purposeful contact exploitation rather than simple avoidance. Effective humanoids must use their bodies to interact with the world as humans do, such as bracing against a wall for balance, blocking objects for safety, or ducking under obstacles. These contact-aware skills are essential for greater autonomy, robustness, and physical intelligence in robots.
Reasoning about contact remains challenging for humanoids [3, 4, 5]. Traditional optimization-based methods [6, 7] struggle with the combinatorial complexity of real-time contact scheduling and are sensitive to model inaccuracies, reducing adaptability to unforeseen situations. Parallelized simulation [8] has enabled on-policy RL to succeed in robot control for quadrupeds [9], bipeds [10], and humanoids [11]. However, these methods are sample-inefficient, especially with visual inputs [9], and struggle with multi-task learning and complex scene interactions.
@@
We address this by integrating a learned world model with sampling-based Model Predictive Control (MPC). Our approach trains a scalable world model from a random, demonstration-free offline dataset, predicting future outcomes in compressed latent space rather than raw pixels, and understanding action consequences. We introduce a surrogate value function to guide planning, allowing MPC to efficiently evaluate candidate action sequences. This synergy enables agile, vision-based contact planning for humanoids across tasks with improved data efficiency and performance.
@@
    Agile and Robust Real-world Visual Contact Planning: We validate the proposed framework on a physical humanoid robot, demonstrating multiple novel agile and robust contact planning tasks solely from ego-centric depth images and proprioceptive feedback.
Figure 2: World Model Training Pipeline. The pipeline begins with the offline data collection process shown in (a), where a dataset $\mathcal{D}$ of trajectories is generated by applying randomly sampled high-level actions (end-effector position $p_{ee}^{\top}$ and body height $h_{body}$) to a simulated humanoid equipped with a trained low-level policy. This dataset is then used to train the world model, as depicted in (b). At each timestep $t$, an observation $o_{t}$, consisting of a depth image and proprioception, is encoded into a stochastic latent state $z_{t}$, which is then decoded to produce a reconstruction $\hat{o}_{t}$. Concurrently, a recurrent network updates its latent state $h_{t}$ based on the previous state and action. From the combined latent state $(h_{t},z_{t})$, the model predicts (i) $\hat{z}_{t}$, a prior sample of the stochastic latent state; (ii) $\hat{d}_{t}$, the termination probability; and (iii) $\hat{Q}_{t}$ a surrogate action-value guiding the planner in evaluating different actions. All of these predictions are optimized against the ground-truth data from the offline trajectories, enabling the model to learn both the environment’s dynamics and a robust value function for planning.
@@ II-B Learning-Based Contact Planning
Learning-based approaches for real-world contact planning have shown remarkable potential, enabling dynamic skills [10, 9, 3, 5, 18, 19]. However, three significant challenges remain largely unaddressed. First, interaction with both dynamic and static objects is limited. Most existing work on legged locomotion based on simplified 2.5D elevation maps [4], which cannot represent dynamic or overhanging obstacles such as a moving ball or an archway. Second, sample efficiency remains a bottleneck. Modern approaches rely heavily on synthetic data [8], but the computational cost of rendering visual input makes on-policy RL algorithms prohibitively expensive [9]. The sparse and discontinuous nature of contact events also poses significant challenges for model-free methods, which often require extensive training or carefully designed inductive biases to guide exploration effectively [5, 20]. Finally, multi-task capability is limited. While current policies can be trained to solve a single task with a specific reward function, they often fail to generalize across interactions with different objects or adapt to variations in task definitions [21].
@@ II-C Planning with Robotic World Models
The world model [22] is a learned, internal model of an environment that allows an robot to predict future outcomes based on actions. The emphasis is not on explicitly predicting the future, but rather on predicting its abstract representation. Using a learned world model for model-based planning offers a path toward better generalization and data efficiency, presenting a promising solution for contact planning [23, 24].
Early work on world models [22] was defined as enabling efficient policy learning in reinforcement learning [23, 24]. Today, the definition has evolved as a generative AI system [25, 26] capable of understanding and simulating the physical world’s causal relationships, dynamics, and interactions. In control and robotics, a similar path is being explored by learning neural dynamics models to represent complex systems [27, 28, 29, 30]. These learned models could be integrated into frameworks such as sampling-based MPC [31, 32, 16] to achieve highly adaptive control [33, 4, 34].
@@ III Methods
This section begins by outlining our hierarchical control framework in Section III-A, composed of a low-level policy and our high-level planner. The subsequent sections then detail the two components of this high-level planner: the world model’s architecture and training (Section III-B), and the value-guided sampling MPC (Section III-C).
@@
Figure 3: Value-Guided Sampling MPC. This figure illustrates how the trained world model is used for planning via value-guided sampling MPC. This process performs open-loop prediction to find the best action sequence starting from a single real observation. At inference time, this process begins by encoding the current observation $o_{t}$ into its latent state $z_{t}$, after which the planner samples a batch of $M=1024$ candidate action sequences over a planning horizon of $N=4$ steps. The world model predicts the future latent state ($h_{t+k},\hat{z}_{t+k}$) by recursively applying its learned dynamics model. At each prediction step, the surrogate value ($\hat{Q}_{t+k}$) evaluates the sampled actions, while the termination signal, $\hat{d}_{t}$, predicts the probability of robot failure, such as falling; if this probability exceeds a threshold of 0.9, all subsequent value estimates, $\hat{Q}$, for that trajectory are set to zero. The planner evaluates $M$ candidate trajectories, where the score for each trajectory is calculated by the objective function $\hat{J}_{N}$ in Eq. (16). This set of scored trajectories is then optimized using the Cross-Entropy Method (CEM) to find the optimal action sequence.
III-A Hierarchical Framework and Data Collection
Our framework consists of two components: a low-level whole-body policy capable of tracking diverse commands, and a high-level planner whose action specifies which commands to track, as shown in Fig. 2.
@@ III-A Hierarchical Framework and Data Collection
The low-level controller tracks commands $c=\left[v^{\top},p_{ee}^{\top},h_{body}\right]^{\top}$, where $v$ is the desired linear velocity, $p_{ee}$ the desired end-effector position, and $h_{body}$ the body height, while maintaining balance. Its observation space is purely proprioceptive, including angular velocity ($\omega$), the projected gravity vector ($g$), the command ($c$), joint positions ($q$), and joint velocities ($\dot{q}$). The controller is represented by a deep neural network and is trained in simulation using PPO, following established approaches [8, 35].
Once a reliable low-level controller is obtained, the high-level planner is designed to account for both contact and task planning. It specifies actions $a_{t}=\left[p_{ee}^{\top},h_{body}\right]^{\top}$ to the low-level controller based on an observation space $o_{t}$ combining proprioception (as in the low-level controller) with visual input from a downsampled 64×48 ego-centric depth image. We exclude base velocity $v$ from the planner’s action space to force the robot to solve contact-rich problems through postural manipulation, rather than learning locomotion-based evasion strategies. The high-level planner consists of a world model and a sampling-based MPC framework. We train the world model from the offline dataset and, at inference time, use it to evaluate $M$ candidate action sequences over an $N$-step horizon and optimize the action $a_{t}$.
@@
Our offline dataset, $\mathcal{D}$,  is generated by collecting trajectories, $\tau$, from the simulated humanoid. At each timestep, the robot receives an observation $o_{t}$, executes a randomly sampled action $a_{t}$, and in return, receives the reward $r_{t}$ and termination signal $d_{t}$ from the environment. These collected transition tuples $\{o_{t},a_{t},r_{t},d_{t}\}$ are then stored in a final trajectory dataset structured as [Batch, Time, Data]. At each step, we sample the finite differences of planner actions $a_{t}=a_{t-1}+\eta\cdot\delta$ from $\delta\sim\mathcal{U}(-1,1)$, where $\eta$ is a scalar step that controls the magnitude of the update. This step is performed after normalizing the task space $\left[p_{ee}^{\top},h_{body}\right]^{\top}$ of the low-level controller. $\eta$ is set to $0.32$. The purpose of using such a method to sample the actions is to (1) avoid using any demonstration, which is expensive to obtain for the whole-body commands of a humanoid, and (2) heuristically avoid ineffective and jitter behavior data far outside the command range.
@@
III-B Ego-Vision Humanoid World Model
Prior auto-regressive models learn system dynamics by mimicking existing controllers for continuous tasks like velocity tracking [29]. However, when applied to high-dimensional image observations, this pixel-prediction approach suffers from compounding errors over long horizons. This issue is exacerbated in contact-aware scenarios where defining a goal trajectory in pixel space is often intractable.
To address this, we draw inspiration from general world models like Dreamer [36] and JEPA [25]. We focus on predicting abstract latent of future observations, enabling the model to capture more fundamental structures within the data. As illustrated in Fig. 2, our world model is composed of several key components detailed below.
First, the world model leverages a recurrent neural network (RNN) to maintain a deterministic dynamics latent state $h_{t}$, which summarizes dynamics information across time. At each step, a stochastic latent state $z_{t}$, which extracts the abstract latent of the current observation, is inferred from the current observation $o_{t}$ and the latent state $h_{t}$. Similar to an autoencoder, the model is trained to reconstruct the observation $o_{t}$ as $\hat{o}_{t}$ after passing it through a latent bottleneck, which compels the latent state $z_{t}$ to encode the most salient and abstract features. For notational simplicity, we let $\phi$ denote the parameters of all world model components, with $q_{\phi}$ and $p_{\phi}$ representing the encoder and decoder, respectively. The overall process can then be expressed as:
@@ III-B Ego-Vision Humanoid World Model
We assume $z_{t}$ and $\hat{o}_{t}$ to follow the Gaussian distribution.
In addition, we introduce a model that estimates the stochastic latent without using the current observation: given $h_{t}$, it yields a latent $\hat{z}_{t}\sim p_{\phi}(\hat{z_{t}}\mid h_{t})$ that closely approximates $z_{t}$, thereby enabling rollouts in latent space.
Different from Dreamer [36], we need to consider an architecture that better addresses the challenges unique to robotics in the real world, such as (1) significant partial observability, (2) high sensor noise, and (3) sparse contact. These factors make it difficult to predict contact-aware rewards from the observation. Therefore, we design specialized heads that, conditioned on the latent state $(h_{t},z_{t})$ and a candidate action $a_{t}$, directly estimate the expected long-term outcome. Specifically, we predict a termination probability $\hat{d}_{t}$ and a surrogate value $\hat{Q}_{t}$, which represents the expected cumulative return. This allows the robot to evaluate the potential response to different actions directly from its learned latent state.
@@
The model is optimized by minimizing the single-step loss as shown below. This total loss is a simple sum of three main components: a reconstruction loss ($\mathcal{L}_{\text{rec}}$), a joint-embedding predictive loss ($\mathcal{L}_{\text{jep}}$), and a Q-loss ($\mathcal{L}_{\hat{Q}}$):
@@
$\mathcal{L}_{\text{rec}}=\mathbb{E}_{\hat{\mathbf{z}}_{t}\sim q_{\phi}(\mathbf{z}_{t}|\mathbf{o}_{t},\mathbf{h}_{t})}\left[\mathcal{L}_{\text{obs}}+\mathcal{L}_{\text{term}}\right].$ (7)
@@
$\mathcal{L}_{\text{jep}}=D_{\text{KL}}\Big(\text{sg}\big(q_{\phi}(\mathbf{z}_{t}|\mathbf{o}_{t},\mathbf{h}_{t})\big)\ \big|\big|\ p_{\phi}(\mathbf{z}_{t}|\mathbf{h}_{t})\Big)$
$+D_{\text{KL}}\Big(q_{\phi}(\mathbf{z}_{t}|\mathbf{o}_{t},\mathbf{h}_{t})\ \big|\big|\ \text{sg}\big(p_{\phi}(\mathbf{z}_{t}|\mathbf{h}_{t})\big)\Big),$ (8)
@@
$\mathcal{L}_{\hat{Q}}=\mathbb{E}_{\mathbf{\tau}\sim\mathbf{D}}\sum_{t}\mathbb{E}_{\mathbf{z}_{t}\sim q_{\phi}(\cdot|\mathbf{o}_{t},\mathbf{h}_{t})}\left[\left({Q}_{\phi}(\mathbf{h}_{t},\mathbf{z}_{t},\mathbf{a}_{t})-Q_{\text{target}}\right)^{2}\right].$ (9)
Figure 5: Sample efficiency comparison: Our method uses an offline dataset collected from random action, while PPO collects data from environments at every iteration. The x-axis represents the number of step transitions used, while the y-axis shows the reward for each task. A greater value on the x-axis indicates a larger amount of data used, and a higher value on the y-axis signifies better performance. While our method utilizes a dataset of at most 1M steps, we continued to train PPO for a greater number of steps to determine when it could achieve comparable performance.
III-C Value-Guided Sampling MPC
In practical applications involving complex robotics and perception, learning a perfect, optimal state-action value function and greedily maximizing it remains a significant challenge, which is further compounded when extending to observation-action value functions. This difficulty stems from two primary sources:

### v2

@@ Abstract
Enabling humanoid robots to exploit physical contact, rather than simply avoid collisions, is crucial for autonomy in unstructured environments. Traditional optimization-based planners struggle with contact complexity, while on-policy reinforcement learning (RL) is sample-inefficient and has limited multi-task ability. We propose a framework combining a learned world model with sampling-based Model Predictive Control (MPC), trained on a demonstration-free offline dataset to predict future outcomes in a compressed latent space. To address sparse contact rewards and sensor noise, the MPC uses a learned surrogate value function for dense, robust planning. Our single, scalable model supports contact-aware tasks, including wall support after perturbation, blocking incoming objects, and traversing height-limited arches, with improved sample efficiency and multi-task capability over on-policy RL. Deployed on a physical humanoid, our system achieves robust, real-time contact planning from proprioception and ego-centric depth images. Code and dataset are available at our website: https://ego-vcp.github.io/.
@@ I Introduction
Humanoids are expected to advance from dynamic locomotion [24, 26] to intelligent interaction in complex, unstructured environments. Achieving this requires purposeful contact exploitation rather than simple avoidance. Effective humanoids must use their bodies to interact with the world as humans do, such as bracing against a wall for balance, blocking objects for safety, or ducking under obstacles. These contact-aware skills are essential for greater autonomy, robustness, and intelligence in robots.
Reasoning about contact remains challenging for humanoids [12, 25, 16]. Traditional optimization-based methods [28, 34] struggle with the complexity of real-time contact scheduling and are sensitive to model inaccuracies, reducing adaptability to unforeseen situations.
Parallelized simulation [20] has enabled on-policy RL to succeed in robot control for quadrupeds [5], bipeds [15], and humanoids [23]. However, these methods are sample-inefficient, especially with visual inputs [5], and struggle with multi-task learning.
@@
We address this by integrating a learned world model with sampling-based Model Predictive Control (MPC). Our approach trains a scalable world model from a random, demonstration-free offline dataset, predicting future outcomes in the compressed latent space rather than raw pixels, and understanding action consequences. We introduce a surrogate value function to guide planning, allowing MPC to evaluate candidate action sequences efficiently. This synergy enables agile, vision-based contact planning for humanoids across tasks with improved sample efficiency and performance.
@@
    Agile and Robust Real-world Visual Contact Planning: We validate the proposed framework on a physical humanoid robot, demonstrating multiple novel agile and robust contact planning capabilities solely from ego-centric depth images and proprioceptive feedback.
Figure 2: World Model Training Pipeline. The pipeline begins with the offline data collection process shown in (a), where a dataset $\mathcal{D}$ of trajectories is generated by applying randomly sampled high-level actions (end-effector position $p_{ee}^{\top}$ and body height $h_{body}$) to a simulated humanoid equipped with a trained low-level policy. This dataset is then used to train the world model, as depicted in (b). At each timestep $t$, an observation $o_{t}$, consisting of a depth image and proprioception, is encoded into an observation latent $z_{t}$, which is then decoded to produce a reconstruction $\hat{o}_{t}$. Concurrently, a recurrent network updates its latent $h_{t}$ based on the previous state and action. The model predicts (i) $\hat{z}_{t}$, a prior sample of the observation latent; (ii) $\hat{d}_{t}$, the termination probability; and (iii) $\hat{Q}_{t}$, a surrogate action-value guiding the planner in evaluating different actions. All of these predictions are optimized against the ground-truth data from the offline trajectories, enabling the model to learn both the environment’s dynamics and a robust value function for planning.
@@ II-B Learning-Based Contact Planning
Learning-based contact planning has demonstrated potential for dynamic skills [15, 5, 12, 16, 19, 17], yet three key challenges persist. First, interaction capabilities are hindered by simplified 2.5D elevation maps [25], which fail to represent dynamic or overhanging obstacles like moving objects or archways. Second, sample efficiency remains a bottleneck; heavy reliance on synthetic data [20] makes vision-based on-policy RL computationally challenging [5], while the sparse, discontinuous nature of contact complicates exploration for model-free methods [16, 37]. Finally, limited multi-task adaptability prevents policies from generalizing across diverse object interactions or varying task definitions [27].
@@ II-C Planning with Robotic World Models
A world model [8] provides a learned internal model of an environment, enabling robots to predict future outcomes within an abstract latent space. Integrating these models into model-based planning offers a promising trajectory for enhancing generalization and sample efficiency in contact planning [11, 9]. While early frameworks focused on accelerating policy learning [11, 9], contemporary world models have evolved into generative systems [2, 3] capable of simulating complex physical causalities and dynamics. Parallel efforts in robotics leverage neural dynamics models to represent intricate systems [21, 38, 13, 14], which can be integrated into sampling-based MPC [33, 22, 36] for adaptive control [18, 25, 35].
@@ III Methods
Section III-A first describes the low-level controller that executes motor control and defines the high-level planning interface used throughout the paper. We then present our data collection procedure for training the planner in Section III-B. Finally, we detail the two core components of our high-level planner: the architecture and training of the world model (Section III-C), and the value-guided sampling MPC framework that utilizes the world model for test-time planning (Section III-D).
@@
Figure 3: Value-Guided Sampling MPC. This figure illustrates how the trained world model is used for planning via value-guided sampling MPC. This process performs open-loop prediction to find the best action sequence starting from a single real observation. At inference time, this process begins by encoding the current observation $o_{t}$ into its latent state $z_{t}$, after which the planner samples a batch of $M=1024$ candidate action sequences over a planning horizon of $N=4$ steps. The world model predicts the future latent state ($h_{t+k},\hat{z}_{t+k}$) by recursively applying its learned dynamics model. At each prediction step, the surrogate value ($\hat{Q}_{t+k}$) evaluates the sampled actions, while the termination signal, $\hat{d}_{t+k}$, predicts the probability of robot failure, such as falling; if this probability exceeds a threshold of 0.9, all subsequent values for that trajectory are set to zero. The planner evaluates $M$ candidate trajectories, where the score for each trajectory is calculated by the objective function $\hat{J}_{N}$ in Eq. (17). This set of scored trajectories is then optimized using the Cross-Entropy Method (CEM) to find the optimal action sequence.
III-A Low-Level Controller
@@
Our framework utilizes a low-level controller capable of tracking diverse locomotion and manipulation commands. The controller tracks the command vector $c=[v^{\top},p_{ee}^{\top},h_{body}]^{\top}$, where $v$, $p_{ee}$, and $h_{body}$ denote the desired linear velocity, end-effector position, and body height, respectively. Its observation space is purely proprioceptive, comprising angular velocity ($\omega$), the projected gravity vector ($g$), the command ($c$), joint positions ($q$), and joint velocities ($\dot{q}$). The controller is trained in simulation via PPO, following established approaches [20, 31].
III-B Data Collection
Once a reliable low-level controller is obtained, we collect an offline object-interaction dataset for the subsequent vision-based world model training. Our offline dataset $\mathcal{D}$ is generated by collecting trajectories $\tau$ in simulation across three object types: a ball, a wall, and an arch. At each timestep, the robot receives an observation $o_{t}$ including a downsampled $64\times 48$ ego-centric depth image and proprioceptive signals, executes a randomly sampled action $a_{t}=\left[p_{ee}^{\top},h_{body}\right]^{\top}$. We exclude the desired linear velocity $v$ from the planner’s action space to force the robot to solve contact-rich problems through postural manipulation. In return, the robot receives the reward $r_{t}$ and the termination signal $d_{t}$ from the environment. These collected transition tuples $\{o_{t},a_{t},r_{t},d_{t}\}$ are then stored in a final trajectory dataset structured as [Batch, Time, Data].
@@ III-B Data Collection
At each step, we sample the finite differences of planner actions $a_{t}=a_{t-1}+\eta\cdot\delta$ from $\delta\sim\mathcal{U}(-1,1)$, where $\eta$ is a scalar step that controls the magnitude of the update. This step is performed after normalizing the task space $\left[p_{ee}^{\top},h_{body}\right]^{\top}$ of the low-level controller. $\eta$ is set to $0.32$. The purpose of using such a method to sample the actions is to (1) avoid using any demonstration, which is expensive to obtain for the whole-body commands of a humanoid, and (2) avoid ineffective and jittery behavior data from random sampling. For a detailed description of the process, see the illustration video on our website and code.
@@
III-C Ego-Vision Humanoid World Model
Prior auto-regressive models learn system dynamics by mimicking existing controllers for continuous tasks such as velocity tracking [13]. However, when applied to high-dimensional image observations, this pixel-prediction approach suffers from compounding errors over long horizons. This issue is exacerbated in contact-aware scenarios where defining a goal trajectory in the pixel space is often intractable.
To address this, we draw inspiration from general world models such as Dreamer [10] and JEPA [2]. We focus on predicting the abstract latent states of future observations, enabling the model to capture more fundamental structures within the data. As illustrated in Fig. 2, our world model is composed of several key components detailed below.
First, the world model leverages a recurrent neural network (RNN) to maintain a deterministic dynamics latent state $h_{t}$. At each step, a stochastic latent state $z_{t}$, which extracts the abstract latent of the current observation, is inferred from the current observation $o_{t}$ and the latent state $h_{t}$. Similar to an autoencoder, the model is trained to reconstruct the observation $o_{t}$ as $\hat{o}_{t}$ after passing it through a bottleneck, which compels the latent state $z_{t}$ to encode the most salient and abstract features. For notational simplicity, we let $\phi$ denote the parameters of all world model components, with $q_{\phi}$ and $p_{\phi}$ representing the encoder and decoder, respectively. The overall process can then be expressed as:
@@ III-C Ego-Vision Humanoid World Model
We model both $z_{t}$ and $\hat{o}_{t}$ as Gaussian distributions.
Furthermore, we introduce a model that estimates the stochastic latent without using the current observation: given $h_{t}$, it predicts a latent $\hat{z}_{t}\sim p_{\phi}(\hat{z}_{t}\mid h_{t})$ that closely approximates $z_{t}$, thereby enabling rollouts in latent space.
Different from Dreamer [10], we need to consider an architecture that better addresses the challenges unique to robotics in the real world, such as (1) significant partial observability, (2) high sensor noise, and (3) sparse contact. These factors make it difficult to predict contact-aware rewards from the observation. Therefore, we design specialized heads that, conditioned on the latent $(h_{t},z_{t})$ and a candidate action $a_{t}$, directly estimate the expected long-term outcome. Specifically, we predict a termination probability $\hat{d}_{t}$ and a surrogate value $\hat{Q}_{t}$, which represents the expected cumulative return. This allows the robot to evaluate the potential response to different actions directly from its learned latent.
@@
The model is optimized by minimizing the loss as shown below. This total loss is a simple sum of three main components: a reconstruction loss ($\mathcal{L}_{\text{rec}}$), a joint-embedding predictive loss ($\mathcal{L}_{\text{jep}}$), and a Q-loss ($\mathcal{L}_{\hat{Q}}$):
@@
$\mathcal{L}_{\text{rec}}=\mathbb{E}_{{{z}}_{t}\sim q_{\phi}({z}_{t}|{h}_{t},{o}_{t})}\left[\mathcal{L}_{\text{obs}}+\mathcal{L}_{\text{term}}\right].$ (7)
@@
$\mathcal{L}_{\text{jep}}=D_{\text{KL}}\Big(\text{sg}\big(q_{\phi}({z}_{t}|{h}_{t},{o}_{t})\big)\ \big|\big|\ p_{\phi}({\hat{z}}_{t}|{h}_{t})\Big)$
$+D_{\text{KL}}\Big(q_{\phi}({z}_{t}|{h}_{t},{o}_{t})\ \big|\big|\ \text{sg}\big(p_{\phi}({\hat{z}}_{t}|{h}_{t})\big)\Big),$ (8)
@@
$\mathcal{L}_{\hat{Q}}=\mathbb{E}_{{\tau}\sim\mathcal{D}}\sum_{t}\mathbb{E}_{{z}_{t}\sim q_{\phi}({z}_{t}|{h}_{t},{o}_{t})}\left[\left({Q}_{\phi}({h}_{t},{z}_{t},{a}_{t})-Q_{\text{target}}\right)^{2}\right].$ (9)
Figure 5: Sample efficiency comparison: Our method uses an offline dataset collected from random actions, while PPO collects data from environments at every iteration. The x-axis represents the number of step transitions used, while the y-axis shows the reward for each task. A greater value on the x-axis indicates a larger amount of data used, and a higher value on the y-axis signifies better performance. While our method utilizes a dataset of at most 1M steps, we continued to train PPO for a greater number of steps to determine when it could achieve comparable performance.
III-D Value-Guided Sampling MPC
In practical applications involving complex robotics and perception, learning a perfect observation-action value function and greedily maximizing it remains a significant challenge. This difficulty stems from two primary sources: